The number of females employees selected if five employees are selected by lottery from a group of six male and six female employees.

1
Expert's answer
2022-03-04T05:19:13-0500

Let "X=" the number of females employees selected. The possible values of the random variable "X" are "0, 1, 2,3,4,5."

There are "\dbinom{6+6}{5}=792"


"P(X=0)=\dfrac{\dbinom{6}{0}\dbinom{6}{5-0}}{\dbinom{12}{5}}"

"=\dfrac{1(6)}{792}=\dfrac{1}{132}"


"P(X=1)=\dfrac{\dbinom{6}{1}\dbinom{6}{5-1}}{\dbinom{12}{5}}"

"=\dfrac{6(15)}{792}=\dfrac{15}{132}"



"P(X=2)=\dfrac{\dbinom{6}{2}\dbinom{6}{5-2}}{\dbinom{12}{5}}"

"=\dfrac{15(20)}{792}=\dfrac{50}{132}"



"P(X=3)=\dfrac{\dbinom{6}{3}\dbinom{6}{5-3}}{\dbinom{12}{5}}"

"=\dfrac{20(15)}{792}=\dfrac{50}{132}"





"P(X=4)=\dfrac{\dbinom{6}{4}\dbinom{6}{5-4}}{\dbinom{12}{5}}"

"=\dfrac{15(6)}{792}=\dfrac{15}{132}"





"P(X=5)=\dfrac{\dbinom{6}{5}\dbinom{6}{5-5}}{\dbinom{12}{5}}"

"=\dfrac{6(1)}{792}=\dfrac{1}{132}"